# Holistic Optimization Framework for FPGA Accelerators

### v4

@@ 6.2. Experimental Evaluation
For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
@@ 6.3. Comparison
Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
@@
2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
@@ 6.4. Scalability
This efficiency stems from Prometheus’ ability to explore a larger optimization space while still maintaining fast solution times. The key reason is that all optimization techniques are seamlessly integrated within the design space. This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity. As a result, even with a broader search space, Prometheus efficiently converges to a solution in 21.37 seconds.
@@ 7. Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024). DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS (Basalama and Cong, 2025) automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower (Amiri et al., 2021) automates FPGA dataflow development but limits parallelism. Frameworks like (Chen et al., 2024; Xiang et al., 2022), built on HeteroCL (Lai et al., 2019), optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays (Wang et al., 2021; Cong and Wang, 2018; Jia et al., 2020; Lai et al., 2020; Basalama et al., 2023) offer efficient computation for specific patterns but lack generalization. Application-specific frameworks (Chi et al., 2018; Hegarty et al., 2014; Natale et al., 2016; de Fine Licht et al., 2021; Singh et al., 2020; Denzler et al., 2023; Korol et al., 2022; Bacis et al., 2017) demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA (Guo et al., 2023a, b) enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
@@
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b). However, these approaches lack effective integration of dataflow optimization techniques.
@@
Code Transformation  Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations (Pouchet et al., 2013b) leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works (Zhao et al., 2022; Zhao and Cheng, 2021) selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper (Pouget et al., 2025b) described in Section 2 generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
@@
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling (Li et al., 2021) focus on CPUs, while Wedler (Shi et al., 2023) optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
@@ 8. Conclusion
Through extensive performance evaluations on computation-bound kernels, we demonstrated that Prometheus outperforms state-of-the-art frameworks.
Furthermore, Prometheus’ automatic design space exploration significantly reduces the manual effort required in FPGA development. By automatically generating HLS-C++ code, OpenCL host code, and FPGA bitstreams, the framework streamlines the deployment process, making FPGA acceleration more accessible to a broader range of applications.
@@ Acknowledgments
This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.

### v5

@@ 6.2. Experimental Evaluation
For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we manually adjust the NLP constraints based on the specific congestion issue and regenerate the HLS-C++ code, a process that typically takes only a few minutes. As our NLP model includes many constraints, both global and per SLR. If congestion occurs, we can address it by adjusting the relevant constraint. For example, if the issue is due to resource overutilization, we can retain the previous NLP solution and simply tighten the resource utilization constraint for the specific SLR. It is worth noting that different regeneration strategies are possible, depending on whether we prioritize achieving a high-quality result (QoR) or obtaining a solution quickly. For now, this decision is left to the user. This step is currently performed manually to better understand the root causes of congestion, but it could be automated through analysis of the compiler-generated log files.
@@ 6.3. Comparison
Sisyphus Prometheus
Kernel GF/s BRAM (%) DSP (%) FF (%) LUT (%) GF/s BRAM (%) DSP (%) FF (%) LUT (%)
Madd 1.96 1 29 22 38 2.71 3 7 10 11
2-madd 3.89 1 58 59 91 8.99 6 14 21 24
3-madd 3.91 2 58 75 117 14.00 9 21 31 37
2mm 195.09 3 73 49 72 308.38 19 71 41 63
3mm 178.97 16 98 86 122 368.36 20 83 52 85
Gemm 227.09 6 51 42 72 419.14 18 53 33 45
Gemver 17.93 24 116 56 68 22.51 31 53 34 54
Mvt 5.54 23 2 6 10 14.65 6 22 16 51
Table 7. RTL Evaluation of Performance and Resource Utilization for Sisyphus and Prometheus
Table 7 presents a comparison between Prometheus and Sisyphus on additional kernels. As shown, Prometheus generally achieves higher throughput and better overall resource utilization, except for BRAM usage, which is higher due to the use of double buffering. Notably, the 3-madd kernel shows a significant performance gain, as it enables concurrent execution of independent tasks. While similar benefits are also observed for 2mm and 3mm, the improvement is less pronounced. Despite achieving a 2.65x speedup on Mvt, we observe that Prometheus consumes more resources than Sisyphus. This is because Sisyphus attempts to unroll the loop with a large unroll factor and apply pipelining, but the compiler fails to achieve an initiation interval (II) of 1 and instead settles for II = 36. As a result, although they lose performance, the larger II significantly reduces the resource usage.
@@
Prometheus achieves a 77.16× performance improvement over AutoDSE, the state-of-the-art (SOTA) model-free approach. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, a recent NLP-based approach, Prometheus demonstrates a notable 2.59× performance boost, leveraging an average of 3.88× more DSPs, 1.04× more BRAM, 1.31× more LUTs, and 1.33× more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
@@
[…]
@@ 6.4. Scalability
This efficiency stems from Prometheus’ ability to explore a larger optimization space while still maintaining fast solution times. The key reason is that all optimization techniques are seamlessly integrated within the design space. This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity.
In the case of 3mm, which has six statements (including initialization), Sisyphus evaluates the product of all possible permutations since it relies on shared buffers. In contrast, Prometheus employs dataflow, which requires preserving the order of data between tasks that communicate through FIFOs. As a result, although each statement could in principle lead to 2! or 3! loop permutations, many of these permutations become invalid under dataflow constraints when tasks must communicate via FIFOs. This substantially reduces the effective search space. Consequently, even with a broader search space, Prometheus converges efficiently to a solution in just 21.37 seconds.
@@ 7. Related Work
Recent years have seen the development of numerous frameworks and tools addressing the challenges of optimizing FPGA accelerators. While many tackle specific subproblems—such as pipelining, tiling, or dataflow—few propose holistic solutions integrating all stages of the transformation and scheduling pipeline. We categorize the most relevant related works according to key technical axes addressed by our framework: dataflow generation, shared-buffer optimization, code transformation, and tiling with padding. We also compare against general-purpose frameworks when applicable.
@@
Dataflow –   Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024). DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS (Basalama and Cong, 2025) automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower (Amiri et al., 2021) automates FPGA dataflow development but limits parallelism. Frameworks like (Chen et al., 2024; Xiang et al., 2022), built on HeteroCL (Lai et al., 2019), optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays (Wang et al., 2021; Cong and Wang, 2018; Jia et al., 2020; Lai et al., 2020; Basalama et al., 2023) offer efficient computation for specific patterns but lack generalization. Application-specific frameworks (Chi et al., 2018; Hegarty et al., 2014; Natale et al., 2016; de Fine Licht et al., 2021; Singh et al., 2020; Denzler et al., 2023; Korol et al., 2022; Bacis et al., 2017) demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA (Guo et al., 2023a, b) enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
@@
Shared Buffer –   Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b). However, these approaches lack effective integration of dataflow optimization techniques.
@@
Code Transformation –   Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations (Pouchet et al., 2013b) leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works (Zhao et al., 2022; Zhao and Cheng, 2021) selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper (Pouget et al., 2025b) described in Section 2 generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
@@
Tiling and Padding –   Tiling is essential for balancing computation and communication. While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling (Li et al., 2021) focus on CPUs, while Wedler (Shi et al., 2023) optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
Discussion –  In summary, while existing works address specific optimization aspects such as dataflow generation, pragma insertion, or tiling, they remain fragmented solutions. Prometheus distinguishes itself by integrating these axes into a unified optimization flow. In particular, it combines dataflow construction, shared-buffer optimization, and code transformation with tiling and padding in a single framework. Furthermore, Prometheus automates the generation of HLS-compatible C++ and OpenCL code and introduces SLR-aware scheduling. This holistic approach, together with public availability on GitHub, provides a significant advance over prior FPGA optimization frameworks.
8. Scopes of Application
Prometheus operates on affine C/C++ programs—a deliberate and powerful design choice. This class of programs captures a vast number of real-world, compute-intensive workloads found in domains such as dense linear algebra, signal processing, image filtering, and many AI kernels. These patterns are not only ubiquitous but also performance-critical in modern FPGA applications, and our evaluation covers typical patterns of computation and data reuse occurring in these computations.
Through MLIR (Lattner et al., 2021) and emerging tools like Allo and Stream-HLS, our framework can seamlessly target programs written in higher-level languages such as Python. By translating these to affine MLIR representations and regenerating C/C++ code (Moses et al., 2021), we enable a broad spectrum of users—from ML practitioners to scientific programmers—to benefit from hardware acceleration without sacrificing optimization quality.
Affine computations can be exactly analyzed at compile-time, at the level of every operation and data element accessed (Feautrier, 1992, 1991), leading to accurate analysis of loop bounds, memory access patterns, data dependencies, and data reuse. This enables precise performance modeling, aggressive design space pruning, and robust application of advanced HLS optimizations such as tiling, unrolling, pipelining, and memory partitioning—tasks that are notoriously difficult in general-purpose compilers or non-affine settings.
Extending to non-affine constructs is possible, but would require significant compromises. For instance, in the presence of indirect accesses like $A[B[i]]$, memory access patterns are no longer statically analyzable, preventing efficient tiling or partitioning. One must either fall back on high-latency off-chip access or resort to over-approximations, e.g., conservatively load entire arrays on-chip—approaches that often degrade parallelism or resource utilization. Our methodology remains applicable, but the design space becomes potentially much more limited due to superfluous constraints arising from over-approximations of the program description.
To achieve full automation and high-performance, our framework relies on the ability to distribute statements in loop nests into distinct loop nests, creating tasks for each generated loop nests afterwards. This property is required for maximal efficiency of our method, and fits typical dense linear algebra kernels for example. But some applications may not expose such property: for example, the time-dependent stencil computations such as Jacobi-2D involve temporal dependencies that prevent full loop distribution, making them less amenable to end-to-end optimization within our current framework. However, the computation performed at each time step can still be optimized effectively using our methodology.
@@ 9. Conclusion
Through extensive performance evaluations on computation-bound kernels, we demonstrated that Prometheus outperforms state-of-the-art frameworks. Furthermore, Prometheus’ automatic design space exploration significantly reduces the manual effort required in FPGA development. By automatically generating HLS-C++ code, OpenCL host code, and FPGA bitstreams, the framework streamlines the deployment process, making FPGA acceleration more accessible to a broader range of applications.
@@ Acknowledgments
This work was supported in part by the NSF award #CCF-2211557, the CDSC industrial partners and the AMD/HACC Program. The authors would like to thank Prof. Luciano Lavagno and Dr. DJ Wang for their helpful discussions and valuable contributions to this research.